### v2

@@
This paper introduces a novel sparse latent factor modeling framework using sparse asymptotic Principal Component Analysis (APCA) to analyze the co-movements of high-dimensional panel data over time. Unlike existing methods based on sparse PCA, which assume sparsity in the loading matrices, our approach posits sparsity in the factor processes while allowing non-sparse loadings. This is motivated by the fact that financial returns typically exhibit universal and non-sparse exposure to market factors. Unlike the commonly used $\ell_{1}$-relaxation in sparse PCA, the proposed sparse APCA employs a truncated power method to estimate the leading sparse factor and a sequential deflation method for multi-factor cases under $\ell_{0}$-constraints. Furthermore, we develop a data-driven approach to identify the sparsity of risk factors over the time horizon using a novel cross-sectional cross-validation method. We establish the consistency of our estimators under mild conditions as both the dimension $N$ and the sample size $T$ grow. Monte Carlo simulations demonstrate that the proposed method performs well in finite samples. Empirically, we apply our method to daily S&P 500 stock returns (2004–2016) and identify nine risk factors influencing the stock market.
@@ 1 Introduction
In the realm of big-data analysis, the study of large-dimensional panel data has gained significant prominence in recent years. Large-dimensional panel data refers to datasets where observations are recorded for multiple individuals or entities over time, resulting in a wealth of information over the spatial and time horizons. Analyzing such datasets is essential for understanding complex economic, financial, and social phenomena. However, as the dimensionality of the data increases, traditional statistical techniques encounter numerous challenges, including multicollinearity, computational complexity, and difficulties in extracting meaningful insights. To address these challenges, latent factor modeling, also referred to as statistical factor modeling in Campbell et al. (1997), has emerged as a powerful and versatile approach in the field of panel data analysis. Latent factor models offer a structured framework for dimension reduction and capturing common sources of variation in high-dimensional data. For example, asset returns in finance are often modeled as functions of a small number of factors; see Ross (1976) and Connor and Korajczyk (1986, 1988). Macroeconomic variables of multiple countries are found to have common components; see Stock and Watson (1989), Gregory and Head (1999), and Forni et al. (2000b). In demand systems, the Engle curves can be expressed in terms of a finite number of factors; see Lewbel (1991). As the dimensions of the panel data systems increase, various factor models have been developed to reduce the dimensionalities under different scenarios. See, for example, the approximate factor models in Chamberlain and Rothschild (1983), Stock and Watson (2002b, a), Bai and Ng (2002), Bai (2003), Fan et al. (2013), Lettau and Pelger (2020), Pelger (2020), and Gao and Tsay (2023, 2024), and the dynamic factor models in Forni et al. (2000a), among others. These models are particularly well-suited for large-dimensional panel datasets where the number of variables (or features) can be moderately large.
@@
One of the most profound challenges in factor modeling is to handle high-dimensional time series data efficiently and effectively. Sparse factor modeling represents a groundbreaking approach to this challenge, offering a powerful technique to extract meaningful insights from complex datasets characterized by a multitude of variables. Sparse factor modeling seeks to identify and capture the underlying structure of data while simultaneously promoting simplicity by selecting only a subset of relevant variables (or features) from the original set. The central idea behind sparse factor modeling is to uncover latent factors that drive the observed data’s variation while enforcing a degree of sparsity, meaning that only a few variables are deemed essential to explain the data’s structure. This approach is particularly valuable in scenarios where the number of variables far exceeds the number of observations, as it not only reduces computational demands but also enhances the interpretability of the results. Examples of articles concerning this approach include the sparse PCA in Jolliffe et al. (2003), Zou et al. (2006), Shen and Huang (2008), and Witten et al. (2009), and the sparse factor analysis with sparse loadings in Kristensen (2017), among others, where all of the aforementioned works are in line with factor analysis with sparse loadings, implying that each factor process is a linear combination of a small subset of the original panel series only.
@@
It is widely known that the key idea in a large-dimensional factor model is that the dimensions of both the timeline and the cross-section of the data are large, and most of the co-movements can be explained by a few factors. These factors and loadings are usually estimated by the conventional principal component analysis (PCA) method due to its ability to parsimoniously capture much of the information in a large number of variables. However, the resulting latent PCs or factors are typically linear combinations of all cross-sectional units/variables, which are usually hard to interpret. The sparse PCA restricts the cardinality of the weight vectors for the PCs, or equivalently, the loadings for the factor processes, so the PCs are sparse linear combinations of the underlying variables. Consequently, the PCs or factors are only linear combinations of a small subset of the cross-sectional units, which facilitates the interpretations of the resulting PCs or factors. Nevertheless, there is little literature concerning the justificaton of the sparsity assumption in the loadings. For example, the assumption may not be appropriate for financial returns, where the exposure of the returns to a market factor is universal and non-sparse, as pointed out in Pelger and Xiong (2022), see also the empirical evidence in Gao and Tsay (2021).
@@
In view of the above discussion, this paper marks a further development in the sparse factor modeling of large-dimensional panel data from a different perspective. Motivated by the empirical success of a general approximate factor model, see Pelger (2020) and the references therein, we also adopt such a similar framework. Unlike the sparse factor modeling approach with sparse loadings in existing literature, we assume the latent factor processes are sparse over time, while the loadings can be nonzero in general settings. Specifically, for a given time period of economic or financial series, we postulate that cross-sectional units may exhibit varying strengths of co-movement over time. During certain periods, the co-movement may not be significant enough, allowing idiosyncratic terms to dominate the co-movement, potentially resulting in the absence of a systematic risk factor. See the empirical evidence of the dominance of idiosyncratic risk in Campbell et al. (2001) and Goyal and Santa-Clara (2003), among others. Consequently, the loading vector of each panel series can be treated as a linear combination of the corresponding series over the entire timeline. When a factor process is sparse, the corresponding loading associated with each series is a linear combination of the series over a small subset of the entire time horizon. This implies that only specific data points contribute to the strength of the loading for each series. The interpretation of this approach is quite straightforward. For instance, asset returns in finance may experience periods of high volatility while also exhibiting relatively stable patterns during certain timeframes. It is known that clear and significant co-movement patterns do emerge during certain periods, such as when the Federal Reserve raises interest rates to combat inflation or during financial crises triggered by currency issues, debt, banking problems, or subprime mortgage defaults. Therefore, our proposed approach can facilitate researchers in economics and finance in discovering the connections between specific policies or events and the co-movements of economic or financial systems.
@@
In this paper, we first introduce approximate factor models and their estimation procedures using the conventional PCA or the asymptotic PCA (APCA) method introduced by Connor and Korajczyk (1986, 1988). We then present a sparse APCA technique by formulating a sparse factor modeling framework with sparse factors over time horizons. Under the $\ell_{0}$-constraint imposed on the factors, we propose a truncated power method to estimate the sparse factors in the one-factor case, corresponding to the hard-thresholding technique in the machine learning literature. Additionally, we develop a sequential deflation estimation procedure for multi-factor cases, utilizing the truncated power method along with certain projection techniques. Furthermore, a data-driven approach is also proposed to identify the sparsity of risk factors over the time horizon using a novel cross-sectional cross-validation method. Theoretical properties of the proposed estimators are established under mild conditions as both the dimension and sample size tend to infinity. Simulated and real data examples are used to illustrate the proposed method. We empirically examine specific events associated with the identified sparse factors that systematically influence the stock market. Our findings reveal nine significant time factors that may systematically affect the financial market. To the best of our knowledge, this is the first study in the literature to consider sparse factors over the timeline, and we establish theoretical guarantees for the estimation procedure. The proposed approach provides an effective way to bridge time and significant risk factors, which play a crucial role in studying the dynamics of large-dimensional panel systems.
@@
Note that the proposed framework not only marks a clear distinction from existing sparse PCA or sparse factor modeling in the literature in terms of formulation and interpretation, but the technique used in this paper also differs from those employed in previous works. For example, in the sparse PCA literature, such as Zou et al. (2006), Shen and Huang (2008), and Witten et al. (2009), the true sparse eigenvectors are assumed to be constant vectors, and regularized estimation algorithms are proposed to estimate these sparse eigenvectors. In the realm of sparse factor modeling, such as in Kristensen (2017) and Uematsu and Yamagata (2022), the loadings are assumed to be sparse and non-random, with $\ell_{1}$-regularized methods used to estimate the sparse loadings. Conversely, in the sparse proximate factor modeling approach of Pelger and Xiong (2022), only the factor weights are assumed to be sparse, while the resulting loading and factor processes are non-sparse. In contrast with the aforementioned frameworks, the proposed sparse APCA method assumes that the random factor processes exhibit a sparse structure, meaning each factor data point can be either a random variable or zero over the timeline. A zero-factor point suggests that the return series exhibits minimal co-movement or that prices remain relatively stable at that particular time. Our estimation procedure is developed under $\ell_{0}$-constraints on the factors. Therefore, our theoretical analysis differs from all the previously mentioned approaches, as the dependence structure along the timeline differs from that across the cross sections. Consequently, our method offers a way to link specific time points with the relevant risk factors that systematically influence the panel of interest.
@@
The contributions of this paper are multi-fold. First, to the best of our knowledge, this is the first study to consider sparse factors over the time horizon, providing economists with a novel approach to bridging the gap between time and risk in economic and financial systems. Second, unlike the $\ell_{1}$-penalized methods in sparse PCA, we propose using a truncated power method based on an $\ell_{0}$-constraint for the one-factor case and a sequential optimization procedure for estimating multiple factors. This extends the asymptotic PCA technique developed by Connor and Korajczyk (1986, 1988) for factor analysis to high dimensions. Third, rather than using the traditional cross-validation method to identify sparsity structure, we introduce a new cross-validation approach by partitioning the spatial dimensions, enabling the estimation of sparsity parameters through this method. Theoretically, we establish the consistency of the proposed estimators as the dimension and sample size approach infinity. Importantly, we provide theoretical guarantees for the proposed cross-sectional cross-validation method, demonstrating its ability to consistently estimate the sparsity structure over the time horizon of factor models. This offers a rigorous approach to linking time and significant systematic risks across the time horizon.
@@
The rest of the paper is organized as follows. Section 2 introduces the sparse factor modeling framework and its estimation procedure. Section 3 presents asymptotic properties of the estimators obtained in Section 2. Section 4 studies the finite-sample performance of the proposed approach via simulation, and 5 illustrates the proposed procedure with an empirical application. Section 6 concludes. All the proofs and derivations for the asymptotic results are relegated to an online Appendix.
@@
Notation: We use the following notation. For a $p\times 1$ vector ${\mathbf{u}}=(u_{1},...,u_{p})^{\prime}$, $\|{\mathbf{u}}\|_{1}=\sum_{i=1}^{p}|u_{i}|$ is the $\ell_{1}$-norm and $\|{\mathbf{u}}\|_{\infty}=\max_{1\leq i\leq p}|u_{i}|$ is the $\ell_{\infty}$-norm. ${\mathbf{I}}_{p}$ denotes the $p\times p$ identity matrix. For a matrix ${\mathbf{H}}$, its Frobenius norm is $\|{\mathbf{H}}\|=[\mathrm{trace}({\mathbf{H}}^{\prime}{\mathbf{H}})]^{1/2}$ and its operator norm is $\|{\mathbf{H}}\|_{2}=\sqrt{\lambda_{\max}({\mathbf{H}}^{\prime}{\mathbf{H}})}$, where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue of a matrix, and $\|{\mathbf{H}}\|_{\min}$ is the square root of the minimum non-zero eigenvalue of ${\mathbf{H}}{\mathbf{H}}^{\prime}$. $|{\mathbf{H}}|$ denotes the absolute value of ${\mathbf{H}}$ elementwisely. The superscript denotes the transpose of a vector or matrix. We also use the notation $a\asymp b$ to denote $a=O(b)$ and $b=O(a)$ or $a$ and $b$ have the same order of stochastic bound when they are random variables.

### v3

@@
Abstract
This paper introduces a novel sparse latent factor modeling framework using sparse asymptotic Principal Component Analysis (APCA) to analyze the co-movements of high-dimensional time series data. Unlike existing methods based on sparse PCA, which assume sparsity in the loading matrices, our approach posits sparsity in the factor processes while allowing non-sparse loadings. This is motivated by the fact that financial returns typically exhibit universal and non-sparse exposure to market factors. The proposed sparse APCA employs a truncated power method to estimate the leading sparse factor and a sequential deflation method for multi-factor cases under $\ell_{0}$-constraints. Furthermore, we develop a data-driven approach to identify the sparsity of risk factors over the time horizon using a novel cross-sectional cross-validation method. We establish the consistency of our estimators under mild conditions for dependent data as both the dimension $N$ and the sample size $T$ grow. Monte Carlo simulations demonstrate that the proposed method performs well in finite samples. Empirically, we apply our method to daily S&P 500 stock returns from 2004 to 2016. Using textual analysis, we investigate specific events linked to the identified sparse factors and uncover nine key risk factors that influence the stock market.
@@ 1 Introduction
In the realm of big-data analysis, the study of large-dimensional panel data has gained significant prominence in recent years. Large-dimensional panel data refers to datasets where observations are recorded for multiple individuals or entities over time, resulting in a wealth of information over the spatial and time horizons. Analyzing such datasets is essential for understanding complex economic, financial, and social phenomena. However, as the dimensionality of the data increases, traditional statistical techniques encounter numerous challenges, including multicollinearity, computational complexity, and difficulties in extracting meaningful insights. To address these challenges, latent factor modeling, also referred to as statistical factor modeling in Campbell et al. (1997), has emerged as a powerful and versatile approach in the field of panel data analysis. Latent factor models offer a structured framework for dimension reduction and capturing common sources of variation in high-dimensional data. For example, asset returns in finance are often modeled as functions of a small number of factors; see Ross (1976) and Connor and Korajczyk (1986, 1988). Macroeconomic variables of multiple countries are found to have common components; see Stock and Watson (1989), Gregory and Head (1999), and Forni et al. (2000b). In demand systems, the Engle curves can be expressed in terms of a finite number of factors; see Lewbel (1991). As the dimensions of the panel data systems increase, various factor models have been developed to reduce the dimensionalities under different scenarios. See, for example, the approximate factor models in Chamberlain and Rothschild (1983), Stock and Watson (2002b, a), Bai and Ng (2002), Bai (2003), Fan et al. (2013), Lettau and Pelger (2020), Pelger (2020), and Gao and Tsay (2023, 2024), and the dynamic factor models in Forni et al. (2000a), among others.
@@
One of the most profound challenges in factor modeling is to handle high-dimensional time series data efficiently and effectively. Sparse factor modeling represents a groundbreaking approach to this challenge, offering a powerful technique to extract meaningful insights from complex datasets characterized by a multitude of variables. Sparse factor modeling seeks to identify and capture the underlying structure of data while simultaneously promoting simplicity by selecting only a subset of relevant variables (or features) from the original set. The central idea behind sparse factor modeling is to uncover latent factors that drive the observed data’s variation while enforcing a degree of sparsity, meaning that only a few variables are deemed essential to explain the data’s structure. This approach is particularly valuable in scenarios where the number of variables far exceeds the number of observations, as it not only reduces computational demands but also enhances the interpretability of the results. Examples of articles concerning this approach include the sparse PCA in Jolliffe et al. (2003), Zou et al. (2006), Shen and Huang (2008), Johnstone and Lu (2009), Witten et al. (2009), and Ma (2013), and the sparse factor analysis with sparse loadings in Kristensen (2017) and Uematsu and Yamagata (2022), among others, where all of the aforementioned works are in line with factor analysis with sparse loadings, implying that each factor process is a linear combination of a small subset of the original panel series only.
@@
It is widely known that the key idea in a large-dimensional factor model is that the dimensions of both the timeline and the cross-section of the data are large, and most of the co-movements can be explained by a few factors. These factors and loadings are usually estimated by the conventional PCA method due to its ability to parsimoniously capture much of the information in a large number of variables. However, the resulting latent PCs or factors are typically linear combinations of all cross-sectional units/variables, which are usually hard to interpret. The sparse PCA restricts the cardinality of the weight vectors for the PCs, or equivalently, the loadings for the factor processes, so the PCs are sparse linear combinations of the underlying variables. Consequently, the PCs or factors are only linear combinations of a small subset of the cross-sectional units, which facilitates the interpretations of the resulting PCs or factors. Nevertheless, there is little literature concerning the justificaton of the sparsity assumption in the loadings. For example, the assumption may not be appropriate for financial returns, where the exposure of the returns to a market factor is universal and non-sparse, as pointed out in Pelger and Xiong (2022). To illustrate, we examine the risk factors of the daily return data of the S&P 500 stocks studied in Pelger (2020), with $N=332$ stocks over $T=3273$ time points. We first apply PCA to the panel, and the estimated factors $f_{t}$ are shown in Figure S.1 of the Supplement. We then regress each return series $x_{i,t}$ on $f_{t}$, and the resulting $p$-values for testing the significance of the regression coefficients are all close to zero, implying that the dependence of the returns on the factors may be non-sparse. Next, we define a sparse factor $f_{t}^{s}$, where $f_{t}^{s}=f_{t}$ if $f_{t}$ is among the 500 largest values in absolute magnitude, and $f_{t}^{s}=0$ otherwise. We find that ${\sum_{t=1}^{3273}(f_{t}^{s})^{2}}/{\sum_{t=1}^{3273}f_{t}^{2}}=79.4\%$, indicating that the top 500 factor points account for 79.4% of the total factor variance over the entire time horizon. Finally, we calculate the $R^{2}$ values for regressing $x_{i,t}$ on $f_{t}$ and $f_{t}^{s}$, respectively. The results are presented in Figure 1. From the figure, we observe that the $R^{2}$ values based on only 500 factor points represent a substantial proportion of those based on the full factor series. The average ratio of $R^{2}$ values from the sparse to the full factor regressions is 77.5%, suggesting that the explanatory power of the sparse factor over time is substantial and deserves further attention.
Figure 1: $R^{2}$ for regression of $x_{i,t}$ on $f_{t}$ and $f_{t}^{s}$, respectively.
@@
In view of the above discussion, this paper marks a further development in the sparse factor modeling of large-dimensional panel data from a different perspective. Motivated by the empirical success of a general approximate factor model, see Pelger (2020) and the references therein, we also adopt such a similar framework. Unlike the sparse factor modeling approach with sparse loadings in existing literature, we assume the latent factor processes are sparse over time, while the loadings can be nonzero in general settings. Specifically, over a given time period, cross-sectional economic or financial series may exhibit varying degrees of co-movement. During certain periods, weak co-movement allows idiosyncratic components to dominate, resulting in relatively lower systematic risk—see, for example, the empirical evidence in Campbell et al. (2001) and Goyal and Santa-Clara (2003). Since the loading vector for each series reflects a linear combination of the underlying factor series, sparsity—where only the most significant factors are retained—implies that loadings are primarily driven by a subset of factors. This formulation is intuitive in financial applications, where asset returns often alternate between periods of high volatility and relative calm. Co-movements tend to intensify during major events such as interest rate hikes by the Federal Reserve or financial crises related to debt, currency, or mortgage markets. Our proposed method offers a useful framework for uncovering connections between major economic or policy events and the evolving co-movement structure of economic or financial systems.
@@
In this paper, we present a sparse APCA technique by formulating a sparse factor modeling framework with sparse factors over time horizons. Under the $\ell_{0}$-constraint imposed on the factors, we propose a truncated power method to estimate the sparse factors in the one-factor case, corresponding to the hard-thresholding technique widely used in the statistics and machine learning literature. As noted in Johnstone and Lu (2009), a key advantage of hard-thresholding over soft-thresholding is that it preserves the magnitude of the retained signals, whereas soft-thresholding may introduce bias into the estimators. While hard-thresholding has been employed in studies such as Johnstone and Lu (2009) and Ma (2013), their focus is primarily on independent and identically distributed (i.i.d.) data with independent/white noises, where the distinction between factors and loadings therein is minimal. From a theoretical perspective, the presence of serial correlation in both the data and the factors poses additional challenges in establishing the convergence rate of the sample covariance matrix and conducting inference on the estimated loadings. In particular, large deviation inequalities such as Bernstein’s inequality cannot be directly applied to time series data. Furthermore, the thresholding values proposed in Johnstone and Lu (2009) and Ma (2013), which rely on the assumption of normality, are not applicable in our setting, as time series data are generally not normally distributed. For multi-factor cases, we develop a sequential deflation estimation procedure that integrates the truncated power method with projection techniques. Additionally, we introduce a data-driven approach to determine the sparsity of risk factors over time using a novel cross-sectional cross-validation method. Theoretical properties of the proposed estimators are established under mild conditions for dependent data as both the dimension and sample size tend to infinity. Simulated examples are used to illustrate the proposed method. We empirically examine specific events associated with the identified sparse factors that systematically influence the stock market. Our findings reveal nine key time factors that may systematically affect the financial market.
@@
The proposed framework is fundamentally different from existing sparse PCA and sparse factor models in both formulation and interpretation, and it also employs a distinct estimation technique. In the sparse PCA literature (e.g., Zou et al. (2006), Shen and Huang (2008), Witten et al. (2009)), the true sparse eigenvectors are treated as constant vectors, and regularized algorithms are used for their estimation. In sparse factor models (e.g., Kristensen (2017), Uematsu and Yamagata (2022)), the factor loadings are assumed to be sparse and non-random, with estimation based on $\ell_{1}$-regularization. The sparse proximate factor model of Pelger and Xiong (2022) assumes sparsity only in factor weights, while the resulting factor and loading processes remain dense. Time-varying factor models, such as Su and Wang (2017), allow for temporal dynamics of the loadings but do not impose sparsity. In contrast, our sparse APCA framework assumes that the factor processes themselves exhibit sparsity—i.e., each factor entry is either an active random variable or zero at a given time. A zero factor value suggests weaker co-movement at that time point. Our method imposes $\ell_{0}$-constraints directly on the factors and is thus theoretically distinct from prior approaches, which focus on cross-sectional sparsity. This temporal sparsity structure enables us to identify specific time points where systematic risk factors significantly influence the panel.
@@
The contributions of this paper are multi-fold. First, to the best of our knowledge, this is the first study to consider sparse factors over the time horizon, providing economists with a novel approach to bridging the gap between time and risk in economic and financial systems. Second, unlike the $\ell_{1}$-penalized methods in most sparse PCA or factor modeling literature, we propose a truncated power method based on an $\ell_{0}$-constraint for the one-factor case, along with a sequential optimization procedure for estimating multiple factors. Our approach differs from the hard-thresholding methods for i.i.d. data in Johnstone and Lu (2009) and Ma (2013) by explicitly accounting for time series data with both cross-sectional and dynamic dependence. Consequently, our method extends the asymptotic PCA technique developed by Connor and Korajczyk (1986, 1988) to high-dimensional factor analysis. Third, instead of relying on traditional cross-validation to determine sparsity structures, we introduce a novel cross-validation approach that partitions the spatial dimensions. Fourth, most thresholding techniques require specifying a threshold value, as seen in Johnstone and Lu (2009) and Ma (2013). In contrast, our method introduces a selection criterion that determines the number of nonzero factor points without requiring pre-specified threshold values. Theoretically, we establish the consistency of the proposed estimators as the dimension and sample size approach infinity. Importantly, we provide statistical guarantees for the proposed cross-sectional cross-validation method, demonstrating its ability to consistently estimate the sparsity structure over the time horizon of factor models.
@@
The rest of the paper is organized as follows. Section 2 introduces the sparse factor modeling framework and its estimation procedure. Section 3 presents asymptotic properties of the estimators obtained in Section 2. Section 4 studies the finite-sample performance of the proposed approach via simulation, and 5 illustrates the proposed procedure with an empirical application. Section 6 concludes. Some illustrative examples of sparse factors over time horizons, additional tables and figures used in the numerical analyses, all theoretical proofs of the main theorems, and a description of the text data extracted from CNN are relegated to an online Supplement.
@@
Notation: We use the following notation. For a $p\times 1$ vector ${\mathbf{u}}=(u_{1},...,u_{p})^{\prime}$, $\|{\mathbf{u}}\|_{1}=\sum_{i=1}^{p}|u_{i}|$ is the $\ell_{1}$-norm and $\|{\mathbf{u}}\|_{\infty}=\max_{1\leq i\leq p}|u_{i}|$ is the $\ell_{\infty}$-norm. ${\mathbf{I}}_{p}$ denotes the $p\times p$ identity matrix. For a matrix $\mathbf{H}$}, its Frobenius norm is $\|{\mathbf{H}}\|=[\mathrm{trace}({\mathbf{H}}^{\prime}{\mathbf{H}})]^{1/2}$ and its operator norm is $\|{\mathbf{H}}\|_{2}=\sqrt{\lambda_{\max}({\mathbf{H}}^{\prime}{\mathbf{H}})}$, where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue of a matrix, and $\|{\mathbf{H}}\|_{\min}$ is the square root of the minimum non-zero eigenvalue of ${\mathbf{H}}{\mathbf{H}}^{\prime}$. $|\mathbf{H}|$ denotes the element-wise absolute value of $\mathbf{H}$. The superscript denotes the transpose of a vector or matrix. We also use the notation $a\asymp b$ to denote $a=O(b)$ and $b=O(a)$ or $a$ and $b$ have the same order of stochastic bound when they are random variables.